OpenStudy (anonymous):

WILL give YOU MEDAL 4(x-b) = x a) b=4/3x b) b=3/4x c) x=4/3b d) x=3/4b those are all fractions, NOT division problems

OpenStudy (anonymous):

4(x-b)=x 4x-4b=x 3x=4b

OpenStudy (anonymous):

so x would equal \[x=4b/3\]

OpenStudy (anonymous):

That's not an answer choice?

OpenStudy (anonymous):

If you were solving for b... 4(x-b)=x 4x-4b=x -4b=-3x b=3x/4

OpenStudy (anonymous):

Hm... Well is the xs and bs of the multiple choice answer like on the bottom of the fraction, or on the right side of the fraction?

OpenStudy (anonymous):

The right side of the fraction

OpenStudy (anonymous):

Okay so c would be an answer

OpenStudy (anonymous):

but b would also be an answer...

OpenStudy (anonymous):

C? Got it ! Could you maybe do some more? ?_?

OpenStudy (anonymous):

so C or B?

OpenStudy (anonymous):

wait, for b, is the x on the right side? or the bottom of the fraction?

OpenStudy (anonymous):

All are on the right of the fractions

OpenStudy (anonymous):

RAINIAR, I'll answer the next one :p

OpenStudy (anonymous):

Wait, is it B or C?

OpenStudy (anonymous):

I'll post the next one...

OpenStudy (anonymous):

Woah.. they're both the same.

OpenStudy (anonymous):

I'm leaning towards C.

OpenStudy (anonymous):

Okay... sorry, I wasn't sure on that one

OpenStudy (anonymous):

That was a weird question, eh?

OpenStudy (anonymous):

Yea, @*Shadow*, answer the next one lol.

OpenStudy (anonymous):

Haha yea it was. I'm gonna go then. C ya

OpenStudy (anonymous):

Cya

OpenStudy (anonymous):

Solve for y, then find the value of y when given x=-2 6x=7-4y a) -12 b) 7/4 c) 19/4 d) 24 THOSE ARE FRACTIONS. N-O-T DIVISION

OpenStudy (anonymous):

So.. I believe the question is asking you to input -2 into the equation for x. If that is the case 6x*-2 would lead the left side to be -12 and then the right side is already simplified Equation: -12 = 7-4y Subtract 7 from each sides leaving the right sides with -12 - 7 = -4y You get left off with -19 = -4y Now, my way to do this.... Take out the Equal sign and have it -19/-4y Change the signs to positive and you end up with 19/4

OpenStudy (anonymous):

Should be right :p

OpenStudy (anonymous):

C? I've got 2 more if you could PRETTY PLEASE answer them???

OpenStudy (anonymous):

Could You?

OpenStudy (anonymous):

@*Shadow* ?

OpenStudy (anonymous):

Solve for h. A = 1 / 2 bh THATS A FRACTION a) h = 2A / b b) 2A = b / h c) h = 1 / 2 Ab d) h = 2Ab THOSE ARE ALL FRACTIONS. NOT DIVISION